# Rethinking Layer Removal: A Hybrid Pruning Framework Combining Layer Removal and Singular Value Selection for Efficient LLM Compression

### v2

@@ A.7 Evaluation Results on LongBench
A.8 Compression Costs of GRASP
To assess the compressed costs of GRASP, we measure its compression time and peak memory consumption when compressing LLaMA2-7B on an NVIDIA A100 GPU under 25% compression ratio. These metrics, summarized in Table 11, indicate that GRASP achieves a good balance between compression time and memory consumption.
Model
Average
Compression Time(h)
Compression Memory
Consumption(MiB)
LaCO 0.05 45046
ASVD 5.5 17632
SliceGPT 0.6 18195
GRASP 0.16 29050
Table 11: Compression time and peak compression memory usage of GRASP and baselines for LLaMA2-7B under 25% compression ratio on a single A100 GPU.
A.9 Robustness of GRASP towards different Representation Evaluation Dataset
In this section, we provide details of the ablation studies conducted to investigate the impact of representation evaluation datasets and the amount of data used for singular value gradient attribution. Specifically, we selected 512 samples from WikiText-2 (Merity et al., 2017) and C4 (Raffel et al., 2020) as calibration data to assess the performance of GRASP when compressing LLaMA3.1-8B-Instruct under 20% compression ratio. Additionally, we selected 64, 128, 256 and 512 samples from WikiText-2 to examine the robustness of GRASP to the change in the number of calibration data. All calibration data were randomly selected from the training splits of the downstream datasets, ensuring no data leakage. As shown in Figure 5, we can observe that GRASP consistently achieves strong performance, indicating that our method is robust to variations in both the representation evaluation dataset and the number of data. Tables 14 and 15 summarize the results of GRASP when compressing LLaMA3.1-8B-Instruct with different representation evaluation datasets (WikiText-2, C4) and varying numbers of calibration data.
A.10 Robustness of GRASP towards Different Pruning Strategy
@@
Figure 5: Performance of GRASP on LLaMA3.1-8B-Instruct under 20% compression using (a) different representation evaluation datasets (WikiText-2, C4) and (b) varying amounts of calibration data from WikiText-2. GRASP demonstrates limited sensitivity to calibration data changes, with final task performance varying within 4%.
Representation Evaluation Dataset WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
@@
Table 14: Zero-shot performance of LLaMA3.1-8B-Instruct compressed by GRASP under 20% compression using 512 samples from WikiText-2 and C4 as representation evaluation datasets.
Representation Evaluation Dataset WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average

### v3

@@ A.6 Evaluation Results on LongBench
A.7 Robustness of GRASP towards different calibration Dataset
In this section, we provide details of the ablation studies conducted to investigate the impact of calibration datasets and the amount of data used for singular value gradient attribution. Specifically, we selected 512 samples from WikiText-2 (Merity et al., 2017) and C4 (Raffel et al., 2020) as calibration data to assess the performance of GRASP when compressing LLaMA3.1-8B-Instruct under 20% compression ratio. Additionally, we selected 64, 128, 256 and 512 samples from WikiText-2 to examine the robustness of GRASP to the change in the number of calibration data. All calibration data were randomly selected from the training splits of the downstream datasets, ensuring no data leakage. As shown in Figure 4, we can observe that GRASP consistently achieves strong performance, indicating that our method is robust to variations in both the calibration dataset and the number of data. Tables 12 and 13 summarize the results of GRASP when compressing LLaMA3.1-8B-Instruct with different calibration datasets (WikiText-2, C4) and varying numbers of calibration data.
A.8 Robustness of GRASP towards Different Pruning Strategy
@@
Calibration Dataset WikiText-2 PTB Openb. ARC_e WinoG. HeSW ARC_c PIQA MathQA Average
@@
Table 12: Zero-shot performance of LLaMA3.1-8B-Instruct compressed by GRASP under 20% compression using 512 samples from WikiText-2 and C4 as calibration datasets.
Calibration Dataset WikiText-2 PTB Openb. ARC_e WinoG. HeSW ARC_c PIQA MathQA Average